the width of a rectangle is one half the length. the perimeter of the rectangle is 90, what are the dimensions of the rectangle?
Answer by KevinSkullxox(4)

Length = x
Width = 1 1/2 x = 3/2x
Perimeter = 90
Perimeter = 2 X length + 2 X width
= 2x + 2 X 3/2x
= 2x + 3x
= 5x
5x = 90
x = 90/5 = 18
Length = 18
Width = 3/2 X 18 = 27